use 4.9t^2 + vot = s

a coin is tossed downward with an initial velocity of 30 meters per second(v0)at an altitude of 402 meters(s). approximetly how long does it take the coin to reach the ground?

Answer by stanbon(57214)

use 4.9t^2 + vot = s
a coin is tossed downward with an initial velocity of 30 meters per second(v0)at an altitude of 402 meters(s). approximately how long does it take the coin to reach the ground?
-------------------------
-4.9t^2-30t = -402 meters
--------------------------------
4.9t^2 + 30t -402 = 0
---------------
Use Quadratic formula:
t = [-30 +- sqrt(900 - 4*4.9*-402)]/(9.8)
---
t = [-30 +- sqrt(8779.2)]/9.8
Positive solution:
t = [-30+93.7]/9.2
t = 6.924 seconds
=======================================================
Cheers,
Stan H.